HELP!
I do not understand how to equate the rest of the table for this equation.
y = x2 - 2x

x | y
-3
-2
-1
0
1

Answer by checkley71(8403)

x | y
-3 15 y=-3*-3-2*-3 or 9+6=15
-2 8 y=-2*-2-2*-2 or 4+4=8
-1 3 y=-1*-1-2*-1 or 1+2=3
0 0 y=0*0-2*0 or 0-0=0
1 -1 y=1*1-2*1 or 1-2=-1